### v1

# SCBench: A KV Cache-Centric Analysis of Long-Context Methods

Abstract
Long-context Large Language Models (LLMs) have enabled numerous downstream applications but also introduced significant challenges related to computational and memory efficiency. To address these challenges, optimizations for long-context inference have been developed, centered around the KV cache. However, existing benchmarks often evaluate in single-request, neglecting the full lifecycle of the KV cache in real-world use. This oversight is particularly critical, as KV cache reuse has become widely adopted in LLMs inference frameworks, such as vLLM and SGLang, as well as by LLM providers, including OpenAI, Microsoft, Google, and Anthropic. To address this gap, we introduce SCBench (SharedContextBench), a comprehensive benchmark for evaluating long-context methods from a KV cache-centric perspective: 1) KV cache generation, 2) KV cache compression, 3) KV cache retrieval, and 4) KV cache loading. Specifically, SCBench uses test examples with shared context, ranging 12 tasks with two shared context modes, covering four categories of long-context capabilities: string retrieval, semantic retrieval, global information, and multi-task. With SCBench, we provide an extensive KV cache-centric analysis of eight categories long-context solutions, including Gated Linear RNNs (Codestal-Mamba), Mamba-Attention hybrids (Jamba-1.5-Mini), and efficient methods such as sparse attention, KV cache dropping, quantization, retrieval, loading, and prompt compression. The evaluation is conducted on six Transformer-based long-context LLMs: Llama-3.1-8B/70B, Qwen2.5-72B/32B, Llama-3-8B-262K, and GLM-4-9B. Our findings show that sub-$O(n)$ memory methods suffer in multi-turn scenarios, while sparse encoding with $O(n)$ memory and sub-$O(n^{2})$ pre-filling computation perform robustly. Dynamic sparsity yields more expressive KV caches than static patterns, and layer-level sparsity in hybrid architectures reduces memory usage with strong performance. Additionally, we identify attention distribution shift issues in long-generation scenarios.
11footnotetext: Work during internship at Microsoft. Corresponding author.
Figure 1: KV Cache lifecycle. Prior benchmarks focus on single-request, while real-world applications reuse KV cache across requests. We propose SCBench and categorize long-context methods into KV Cache Generation, Compression, Retrieval, and Loading from a KV-cache-centric perspective.
1 Introduction
(a) Two Shared Context Modes
(b) Overview of SCBench
Figure 2: Long-context tasks often involve contexts sharing, e.g., multi-turn dialogues, multi-step reasoning, and repository-level tasks. (a) Illustration of two common shared-context patterns. (b) Overview of tasks and scenarios covered by our benchmark, encompassing four categories of long-context abilities and two shared-context modes.
Long-context capability is becoming a standard for Large Language Models (LLMs), with many of them supporting context windows ranging from 128K to 10M tokens (Reid et al., 2024; Lieber et al., 2024; Dubey et al., 2024; Gradient, 2024). These extended context windows unlock a wide range of real-world applications, such as repository-level code understanding and debugging (Bairi et al., 2024; Park et al., 2023; Liu et al., 2024c; Jimenez et al., 2024), long-document question-answering (Caciularu et al., 2023; Li et al., 2024b), many-shot in-context learning (Agarwal et al., 2024), and long-generation Chain-of-Thought (CoT) reasoning (OpenAI, 2024a; Snell et al., 2024).
Despite the benefit, Long-context inputs also present unique challenges for LLM inference due to high computational costs and memory demands. This has led to the development of efficient long-context solutions leveraging sparsity in various stage of KV cache. In this paper, we introduce an unified analysis framework for efficient long-context methods in a KV cache centric perspective, consisting of the four essential stages of KV cache: 1) KV cache generation, 2) KV cache compression, 3) KV cache retrieval, and 4) KV cache loading. First, KV cache generation, aka prefilling, processes the input prompt and produce the KV cache to be used in decoding. In this stage, sparse attention methods are proposed to reduce the complexity of the attention operation (Child et al., 2019; Beltagy et al., 2020; Jiang et al., 2024). Second, KV cache compression techniques prune KV states to reduce the memory costs in decoding (Xiao et al., 2024b; Li et al., 2024c). Third, KV cache retrieval aims to skip KV cache generation of an incoming request, and instead retrieves and reuses KV cache from history KV cache pool for more efficient inference (Zheng et al., 2024; Yao et al., 2024a). At last, KV cache loading aims to load only partial of the KV cache for each decoding step to save the memory and computing cost (Tang et al., 2024; Liu et al., 2024b).
However, these methods are only evaluated on single-request benchmarks (Hsieh et al., 2024; Zhang et al., 2024a; Kamradt, 2023; Li et al., 2024a), which fail to covers the full lifecycle of KV cache in real applications. Typically, real-world applications often require reusing prompt memory (i.e., KV cache) and involving multiple requests or multi-round interactions (Qin et al., 2024). The reuse of KV cache, known as prefix caching, is already a crucial component in popular inference frameworks (Zheng et al., 2024; vLLM, 2024) and used by LLM providers (Gemini, 2024; Claude, 2024; OpenAI, 2024b; Azure, 2024). In addition, testing with multiple requests is especially crucial for the long-context methods mentioned earlier, as many achieve efficiency through query-conditioned compression. For instance, Arora et al. (2024) reports that Mamba’s compression of previous information based on the current query can prevent it from answering follow-up queries.
To address this gap, we introduce SCBench, a benchmark designed to evaluate efficient long-context methods that covers the full lifecycle of KV cache in real-world scenarios, particularly for shared context and multi-round interactions where KV Cache is reused for follow-up queries. As shown in Fig. 2b, SCBench assesses four key long-context abilities across 12 tasks with two shared context modes. Each test example includes a shared context and multiple follow-up queries. The four long-context capabilities and their corresponding tasks are:
(a) Performance Across Different Requests
(b) Performance in Different Abilities
Figure 3: Overview of performance results for SCBench. (a) Performance trends of various long-context methods across multiple requests. Methods with $O(n)$ memory cost in decoding show improving performance as requests increase. In contrast, methods with sub-$O(n)$ KV cache in decoding, like KV cache dropping methods, perform well only in the first request. (b) Specific performance of different long-context methods across various long-context capability tasks. All evaluated long-context methods exhibit some loss in Retrieval capability while largely maintaining Global Information processing capability.
Figure 4: Performance of various long-context methods at different compression rates on SCBench using Llama-3.1-8B (Dubey et al., 2024).
  1. 1.
    String Retrieval Capability: A fundamental requirement for long-context LLMs is retrieving relevant context with exact matches from long inputs. We extend previous retrieval tasks like NIAH and Multi-NIAH (Kamradt, 2023; Hsieh et al., 2024) by introducing three comprehensive string retrieval tasks: key-value retrieval, prefix-suffix retrieval, and multi-hop retrieval, measuring capability at different levels of granularity.
  2. 2.
    Semantic Retrieval Capability: Real-world applications often require long-context LLMs to understand semantic meaning before succeeding in retrieval. We considered various semantic retrieval scenarios across different domains, building four distinct tests: RepoQA (Liu et al., 2024c) and long-form QA (covering English, Chinese, and multiple-choice questions) (Zhang et al., 2024a).
  3. 3.
    Global Information Capability: We also assess the capability of long-context LLMs to process and aggregate global information through three tasks: many-shot in-context learning (Agarwal et al., 2024), summarization, and long array statistics (Zhang et al., 2024a).
  4. 4.
    Multi-tasking Capability: In real applications, LLMs often handle multiple tasks with a shared long-context input. Our benchmark evaluates this capability through two tasks: RepoQA with NIAH and summarization with KV retrieval.
In addition, as shown in Fig. 2a, our benchmark includes two typical shared context modes: Multi-turn Mode, where the context is cached within a single session, and Multi-request Mode, where it is cached across multiple sessions.
With SCBench, we conduct an extensive KV cache-centric analysis in the four stages as shown in Fig. 1 (details in Sec. 2). Specifically, we evaluate 13 long-context methods across four stages and eight categories on eight open-source long-context LLMs, including Llama-3.1-8B/70B (Dubey et al., 2024), Qwen2.5-72B/32B (Team, 2024), Llama-3-8B-262K (Gradient, 2024), GLM-4-9B-1M (GLM et al., 2024), Codestal Mamba (team, 2024), and Jamba-1.5-mini (Lieber et al., 2024). These methods span gated linear RNNs (e.g., Codestal Mamba), hybrid models (e.g., Jamba-1.5), sparse attention (e.g., A-shape, Tri-shape, MInference (Jiang et al., 2024)), prompt compression (e.g., LLMLingua-2 (Pan et al., 2024)), KV cache dropping (e.g., StreamingLLM (Xiao et al., 2024b), SnapKV (Li et al., 2024c)), KV cache quantitation (e.g., KIVI (Liu et al., 2024e)), semantic retrieval (e.g., CacheBlend (Yao et al., 2024a)), and KV cache loading (e.g., Quest (Tang et al., 2024), RetrievalAttention (Liu et al., 2024b)), as detailed in Table 1. Additionally, we introduce Tri-shape, a novel, training-free sparse attention method that demonstrates improved first-turn performance in our evaluations.
Our experimental results reveal the following insights: 1) Sub-$O(n)$ memory is almost infeasible in multi-turn decoding, as shown in Fig. 3. Sparse decoding methods (sub-$O(n)$ memory) perform well on the first query but lose accuracy in subsequent requests. In contrast, sparse encoding methods ($O(n)$ memory with $O(n^{2})$computation during pre-filling) can approximate full attention accuracy across multiple queries. 2) Task performance shows varying decline trends, as illustrated in Fig. 3b. Sparse KV cache methods excel in tasks requiring global information, whereas $O(n)$ memory is essential for tasks involving exact match retrieval. 3) All long-context methods experience performance degradation as the budget decreases, as shown in Fig. 4. However, sub-$O(n)$ memory methods exhibit a significant performance drop at a 1/4 compression rate. Methods such as RetrievalAttention and KIVI, which maintain $O(n)$ memory with sparse decoding, sustain higher performance even at higher compression rates. 4) Long-generation scenarios exhibit distribution shift issues, as generation length and the number of rounds increase, the importance distribution of the KV cache changes significantly. This out-of-distribution (OOD) issue leads to performance degradation, even for $O(n)$ memory methods like RetrievalAttention, as observed in extended tasks, as shown in Fig. 3.
Our contributions are as follows:
  • We propose a new benchmark, SCBench, to evaluate long-context methods on multi-round and multi-request scenarios in two typical KV cache reuse scenarios, providing a more realistic assessment.
  • We design an extensive set of downstream tasks, covering four long-context capabilities across 12 subtasks in various domains.
  • We systematically categorize long-context methods from a KV-cache-centric perspective and evaluate 13 different long-context methods (including our newly proposed sparse attention method, Tri-shape) on eight state-of-the-art open-source long-context LLMs using SCBench. Our comprehensive analysis reveals key insights into the effects of sparsity in encoding and decoding, task complexity, and more.
2 A KV Cache-Centric Perspective on Long-Context Methods
Table 1: We evaluated long-context methods on SCBench, where $n$ represents the token size of the input prompt and $m$ represents the generation token size, with $n\gg m$.
Methods Taxonomy Stage P-stage Efficient D-stage Efficient KV Cache Size Prefilling Complexity Decoding Complexity
Codestral Mamba (team, 2024) Gated Linear RNN $O(k)$ $O(kn)$ $O(km)$
Jamba (Lieber et al., 2024) Gated Linear RNN + Full Attention $O(n)$ $O(n^{2})$ $O(nm)$
LLMLingua-2 (Pan et al., 2024) Prompt Compression $O(\alpha n)$ $O(\alpha^{2}n^{2})$ $O(\alpha nm)$
A-shape (Xiao et al., 2024b) Sparse Attention $O(n)$ $O(kn)$ $O(nm)$
Tri-shape
MInference (Jiang et al., 2024)
StreamingLLM (Xiao et al., 2024b) KV Cache Dropping $O(k)$ $O(n^{2})$ $O(km)$
SnapKV (Li et al., 2024c)
PyramidKV (Cai et al., 2024)
KIVI (Liu et al., 2024e) KV Cache Quantitation $O(n)$ $O(n^{2})$ $O(nm)$
CacheBlend (Yao et al., 2024a) KV Cache Retrieval $O(n)$ $O(n^{2})$ $O(nm)$
Quest (Tang et al., 2024) KV Cache Loading $O(n)$ $O(n^{2})$ $O(km)$
RetrievalAttention (Liu et al., 2024b)
Recently, a series of works (Gu & Dao, 2024; Xiao et al., 2024b; Jiang et al., 2024) have explored various strategies to reduce the inference cost of long-context LLMs, enabling their application (Jimenez et al., 2024; OpenAI, 2024a) to downstream tasks at a lower computational expense. In long-context LLM inference, the KV cache plays a pivotal role by effectively reducing the computational overhead during the decoding phase. This importance has led to the development of numerous system-level optimizations (Sheng et al., 2023; Qin et al., 2024) focused on KV cache management and scheduling.
In this work, we propose a novel perspective: these long-context methods can be viewed as optimizations centered around the KV cache at different stages. Specifically, we introduce a KV-cache-centric framework that systematically categorizes long-context methods into four stages: KV Cache Generation, Compression, Retrieval, and Loading, as illustrated in Fig. 1.
Specifically, the four stages of the KV-cache-centric framework are defined as follows:
  1. 1.
    KV Cache Generation: This stage optimizes the efficient generation of KV cache during inference. Techniques include sparse attention (e.g., A-shape, Tri-shape, MInference (Jiang et al., 2024)), SSM or hybrid approaches (e.g., Mamba (Gu & Dao, 2024; Dao & Gu, 2024; Lieber et al., 2024)), and prompt compression (e.g., LLMLingua-2 (Pan et al., 2024)).
  2. 2.
    KV Cache Compression: After generation, the KV cache is compressed before being stored. Methods include KV cache dropping (e.g., StreamingLLM (Xiao et al., 2024b), SnapKV (Li et al., 2024c)) and KV cache quantization (e.g., KIVI (Liu et al., 2024e)).
  3. 3.
    KV Cache Retrieval: Relevant KV cache blocks are retrieved from a storage pool based on the request’s prefix, reducing time-to-first-token (TTFT). Approaches include semantic retrieval methods like CacheBlend (Yao et al., 2024a).
  4. 4.
    KV Cache Loading: This stage dynamically loads the KV cache and computes sparse attention, from KV cache storage (e.g., VRAM, DRAM, SSD, or RDMA) to GPU on-chip SRAM, including Quest (Tang et al., 2024), RetrievalAttention (Liu et al., 2024b), and MagicPIG (Chen et al., 2024).
In our work, we evaluate all four stages of 13 long-context methods, as shown in Table 1. Additionally, we list the KV cache size, pre-filling stage complexity, decoding stage complexity, and whether efficient operations are performed during the pre-filling and decoding stages for each method.
Figure 5: The sparse attention methods framework.
Tri-shape Sparse Attention:
We also introduce a novel training-free sparse attention method, Tri-shape, with improved first-turn accuracy, as shown in Fig. 5. Specifically, in addition to retaining the sink token and local window regions preserved by A-shape, Tri-shape also retains the last window query region, forming a triangular pattern for sparse attention during the pre-filling stage. The motivation arises from the observation from our SCBench that A-shape with dense decoding exhibits significant performance improvement after multiple requests. Tri-shape can notably enhance performance in both turn-0 and multi-request scenarios, as detailed in Sec. 4. Furthermore, it preserves the ability of LLMs to follow instructions, as demonstrated in the case study in §F. This is an example of how our SCBench can provide insights for novel algorithm innovation. Notably, some recent concurrent works (Acharya et al., 2024) have also proposed similar patterns to accelerate the long-context pre-filling stage.
3 Benchmark Building
SCBench comprises 12 tasks covering four long-context abilities: string retrieval, semantic retrieval, global information processing, and multi-tasking, across two shared context modes—multi-turn and multi-request. These tasks span various domains, including code, retrieval, question answering, summarization, in-context learning, multi-hop tracing, and multi-tasking, as shown in Fig. 2b. In total, SCBench includes 931 multi-turn sessions with 4,853 queries, averaging 5 turns per session. Task statistics are provided in Table 2, with examples and configurations in Table 3. Below, we detail the construction of our benchmark.
Table 2: Overview of SCBench tasks.
Task Description Capability Avg. Input Length Avg. Output Length #Sessions / #Turns
Retr.KV Key-value retrieval from many key-value pairs String Retrieval 125K 943 100/500
Retr.Prefix-Suffix Find string with specific prefix and suffix in a dict String Retrieval 112K 914 100/500
Retr.MultiHop Tracking variables assignment in a long input String Retrieval 124K 410 90/450
Code.RepoQA Functions retrieval from a GitHub repo Semantic Retrieval 65K 6,058 88/440
En.QA English Question Answering Semantic Retrieval 198K 272 69/351
Zh.QA Chinese Question Answering Semantic Retrieval 1.5M 322 35/189
En.MultiChoice English Multi-Choice Questions Semantic Retrieval 188K 215 58/299
Math.Find Math computation tasks within long sequence arrays Global Information 120K 172 100/240
ICL.ManyShot Hundreds-shot in-context learning Global Information 22K 975 54/270
En.Sum Summarize a doc given multiple docs as input Global Information 104K 1,170 79/350
Mix.Sum+NIAH Multi-tasking of En.Sum and Needle in A Haystack Multi-tasking 105K 3,441 70/560
Mix.RepoQA+KV Multi-tasking of RepoQA and KV retrieval Multi-tasking 68K 5,318 88/704
Total - - 227K 1,684 931/4,853
3.1 Long-context Task Details
String Retrieval
The most fundamental requirement for long-context LLMs is the ability to identify and retrieve information relevant to a specific query from a lengthy, potentially noisy input. To evaluate this, string retrieval tasks are widely used, where models must retrieve a specific string based on given conditions (Hsieh et al., 2024; Zhang et al., 2024a). Our benchmark incorporates complexity analysis, similar to approaches used in algorithmic problem-solving, such as LeetCode, to design three distinct tasks with varying levels of difficulty. Additionally, by varying the position of the target string, our benchmark further evaluates how well models utilize the full extent of their claimed context window (Kamradt, 2023).
(i) Retrieve.KV: Given a large JSON object containing numerous key-value pairs, the models must accurately retrieve the value corresponding to a specified key (Liu et al., 2024d). The random KVs in this task present significant challenges for long-context LLMs, as the input is often incompressible, requiring strict $O(n)$ space to store. This makes it particularly useful for testing the fuzziness of memory in long-context methods, especially in KV Cache usage. In each session, five KV pairs are retrieved, with the target KVs evenly distributed across the full length of the input.
(ii) Retrieve.Prefix-Suffix: Given a large list of variable-length strings, the models must accurately retrieve a string with a specific prefix and suffix. This task is particularly challenging, as the models need to implement complex functions to match both the prefix and suffix (similar to a prefix tree, with a computational cost of $O(\sum{w_{i}}^{2})$, where $w_{i}$ represents the length of the $i$-th string [^1] https://leetcode.com/problems/prefix-and-suffix-search/). The presence of distractors that share either the prefix or suffix, but not both, prevents models from relying on simple lookup mechanisms or induction heads (Olsson et al., 2022) to solve the task effectively.
(iii) Retrieve.MultiHop: This task, first proposed in RULER (Hsieh et al., 2024), is designed to evaluate the multi-hop tracing capabilities of LLMs within a long input prompt. It requires models to capture and memorize changes in key information from the input context, making it ideal for testing long-context methods in KV cache reuse. Five multi-hop variable assignment chains are embedded throughout the context, and each turn in the test session requires the models to retrieve the exact multi-hop chain, i.e., all variables assigned to a specific value.
Table 3: Task examples and configurations in SCBench. We use different colors to highlight the questions, answers, and distractors in our examples.
Task Source Configuration Example
Retr.KV
Lost in the Middle
(Liu et al., 2024d)
num kv pairs = 2500
len of key & value = 36
metric = Accuracy
Input: {<key #1>: <value #1>, …, <key #100>: <value #100>} Turn 1: The value of the <key #1> is? Answer 1: …<value #1> Turn 2: The value of the <key #20> is? Answer 2: …<value #20>… Turn 3: The value of the <key #40> is? Answer 3: …<value #40>…
Retr.Prefix-Suffix
Ours
size of dict = 6000
len of string = [65, 123)
metric = Accuracy
Input: Dictionary = [<str #1>, <str #2>, …, <str #100>] Turn 1: Prefix: <px #1>; Suffix: <sx #1>. The word with both prefix and suffix from the dict is? Answer: <str> Turn 2: Prefix: <px #2>; Suffix: <sx #2>. Answer: <str>
Retr.MultiHop
RULER
(Hsieh et al., 2024)
num chains = 2
num hops = 2
metric = Accuracy
Input: VAR X1 = 12345 …… VAR Y1 = 54321 …..<noise> VAR X2 = X1 …… VAR Y2 = Y1 ……<noise> VAR X3 = X2 …… VAR Y3 = Y2 ……<noise> Turn 1: Variables that are assigned to 12345? Answer 1: X1 X2 X3 Turn 2: Variables that are assigned to 54321? Answer 1: Y1 Y2 Y3
Code.RepoQA
RepoQA
(Liu et al., 2024c)
func description from GPT-4
metric = Pass@1
Input: <func 1> + <func 2> + … + <func 100> Turn 1: <description of func 1>. Answer 1: <func 1> Turn 2: <description of func 20>. Answer 2: <func 20>
En.QA
Zh.QA
InfiniteBench
(Zhang et al., 2024a)
ground_truth from human
metric = Accuracy
Input: Read the book below and answer a question. <context> Turn 1: <question> Be very concise. Answer 1: …<ans> Turn 2: <question> Be very concise. Answer 2: …<ans>…
En.MultiChoice
InfiniteBench
(Zhang et al., 2024a)
ground_truth from human
metric = Accuracy
Input: Read the book and answer the question. <context> Turn 1: <question> + <Option A,B,C,D>. Answer 1: …<ans> Turn 2: <question> + <Option A,B,C,D>. Answer 2: …<ans>…
Math.Find
Ours
len_array=30000
num_digits=3
metric = Accuracy
Input: <a large array of number> Turn 1: The max number in the array is? Answer 1: …<max number> Turn 2: The max number in the array is? Answer 2: …<max number>
ICL.ManyShot
ManyShotICL
(Srivastava et al., 2023)
num_examples = ~150
Tasks = date, salient, tracking7
metric = Accuracy
Input: ICL Demo. 1 + Demo. 2 + ….. + Demo. 1000 Turn 1: <question>. Answer 1: …<ans> Turn 2: <question>. Answer 2: …<ans>…
En.Sum
Ours
Concatenated arXiv papers
ground_truth from GPT-4
num document = ~8
metric = ROUGE
Input: Doc 1 + Doc 2 + Doc 3 + … + Doc 10. Turn 1: Please summarize Doc 1. Answer 1: … <summary of Doc 1>… Turn 2: Please summarize Doc 3. Answer 2: … <summary of Doc 3>… Turn 3: Please summarize Doc 5. Answer 2: … <summary of Doc 5>…
Mix.Sum+NIAH
Ours
num needle = 5
num document = ~8
metric = ROUGE + Acc
Input: Doc 1 + <Passkeys> + Doc 2 + … + <Passkeys> + Doc 10. Turn 1: Please summarize Doc 1. Answer 1: …<summary of Doc 1>… Turn 2: What is the needle? Answer 2: ..<needle>…
Mix.RepoQA+KV
Ours
num KV pairs = ~100
metric = Pass@1 + Acc
Input: <func 1> + KV pairs + <func 2> + … + KV pairs + <func 100> Turn 1: <description of func 1>. Answer 1: <func 1> Turn 2: The value of the <key #1> is? Answer 2: …<value #1>..
Semantic Retrieval
In addition to string retrieval, many real-world long-context applications require semantic understanding beyond simple string matching, such as retrieving a function based on textual descriptions or answering questions from a long document. These tasks are crucial in SCBench, as lossy long-context methods may struggle to abstract or comprehend information in multi-request scenarios.
(i) Code.RepoQA: This task requires the model to retrieve a specific function (including the function name, input parameters, and full implementation) from a long chunk of source code based on a precise natural language description. Unlike the original RepoQA benchmark (Liu et al., 2024c), our inputs are extended to 64K tokens, with target functions evenly selected based on their position in the codebase. The function descriptions were generated using GPT-4 based on the functions themselves. Additionally, we expanded the range of repositories and programming languages in our test to include Python, C++, Java, PHP, Rust, Go, and TypeScript, compared to the original RepoQA. Each test session involves a GitHub repository, and the model is required to retrieve one function per turn, with a total of 5 turns per session.
(ii) En.QA, Zh.QA, En.MultiChoice: These three tasks are extended from InfiniteBench (Zhang et al., 2024a), which provides high-quality, human-annotated QA tests based on fictional novels to eliminate the influence of external knowledge. These tasks require models to locate and process information within lengthy inputs, performing reasoning through aggregation or filtering to derive answers. There are two primary types of questions: 1) Aggregation involves compiling scattered information throughout the input. An example question is, “How much money in total did A spend on food?” 2) Filtering equires identifying specific information from a larger set. An example question is, “What color dress did A wear when A met B for the second time?” In SCBench, we combine QA pairs that share the same input context to create shared context test sessions.
Global Information Processing
In addition to retrieval, some long-context tasks require leveraging and aggregating global context information, such as summarization, statistical tasks, and in-context learning (Yu et al., 2020; Srivastava et al., 2023; Hao et al., 2022). Our benchmark includes three relevant tasks to assess how well different long-context methods handle global information in multi-request settings.
(i) Many-shot ICL: We use datasets from Big-Bench Hard (Srivastava et al., 2023) to evaluate many-shot in-context learning (ICL) capabilities. This includes three sub-tasks: date understanding, salient error translation detection, and tracking seven shuffled objects. We construct many-shot ICL contexts shared across different turns within a test session. All three sub-tasks are presented as multiple-choice questions with four options provided.
(ii) Math.Find: We extended the math find task from InfiniteBench (Zhang et al., 2024a), expanding from finding only maximum value to multiple statistical values. Given a large array, LLMs are required to find the minimum or median values. LLMs must effectively comprehend global long-context information, perform comparisons, and carry out statistical operations to answer the questions.
(iii) En.Sum: This task uses concatenated academic papers from arXiv as input, with document lengths ranging from 8K to 20K tokens. Ground truth summaries were generated using GPT-4, which was prompted to produce concise one-sentence summaries for each document. The average length of the ground truth summaries is 654 tokens. The target documents for each turn are evenly distributed across the full context length.
Multi-Tasking
In real-world applications, LLMs often handle multiple tasks within a single session using a shared input context. For instance, users might request both summarization and content retrieval simultaneously. To reflect this, we include two multi-tasking tasks in SCBench:
(i) Mix.Sum+NIAH: This task combines document summarization with the Needle in a Haystack (Kamradt, 2023) task using a shared input prompt. A random "needle" is evenly inserted into the En.Sum task’s input (concatenated academic papers). The model alternates between summarization and NIAH retrieval in each test session.
(ii) Mix.RepoQA+KV: This task combines the RepoQA task with KV retrieval using a shared input prompt. Multiple KV pairs are evenly inserted into the RepoQA input (a long chunk of source code). A total of 100 KV pairs are included, with four target KVs and the rest as distractors. The model alternates between RepoQA and KV retrieval in each test session.
3.2 Long-Context Shared Context Modes Details
In addition to the carefully designed long-context tasks, we include two shared context modes to more accurately reflect real-world long-context applications: multi-turn mode and multi-request mode, as shown in Fig. 2b.
(i) Multi-turn Mode: A typical scenario in long-context applications, including long-context chat, multi-step reasoning (e.g., Tree-of-Thought (Yao et al., 2024b)), and long-generation CoT. This mode is relevant to long-context methods with KV cache reuse, as the focus in each turn may shift significantly, potentially causing the models to lose information stored in KV cache. Following Zheng et al. (2023a); Wang et al. (2024), we use ground-truth answers instead of model-generated content as the context for follow-up turns.
(ii) Multi-request Mode: Context sharing can occur across sessions or even users, such as multiple users working on the same code repository. In this case, models can encode the shared context and share the memory (KV cache) across multiple requests. Testing long-context methods in such scenarios is crucial, as some require the query for sparse encoding/decoding. For instance, MInference and SnapKV use the final part of the input (often the query) to estimate the overall sparse pattern. This mode tests how well these methods generalize without having the query.
4 Experiments & Results
Models & Implementation Details
We selected six open-source long-context LLMs for our study: Llama-3.1-8B/70B (Dubey et al., 2024), Qwen2.5-72B/32B (Team, 2024), Llama-3-8B-262K (Gradient, 2024), and GLM-4-9B-1M (GLM et al., 2024), along with two gated linear models: Codestal Mamba 7B (team, 2024), and Jamba-1.5-Mini (Lieber et al., 2024). This selection encompasses Transformer, SSMs, and SSM-Attention Hybrid models, representing some of the most effective context lengths among open-source Long-context LLMs. To ensure result stability, all experiments were conducted using greedy decoding in BFloat16 on four NVIDIA A100 GPUs. We evaluated all models using the HuggingFace or vLLM framework with FlashAttention-2 (Dao, 2024) implementation. Additionally, we employed MInference’s implementation (Jiang et al., 2024) to reduce GPU memory overhead. More information of these models and our infrastructure can be found at §C.1.
Long-Context Method Details
We evaluated eight categories of long-context solutions on our benchmark, including Gated Linear RNNs (e.g., Codestral-Mamba), SSM-Attention hybrid models (e.g., Jamba), sparse attention, KV cache dropping, prompt compression, KV cache quantization, KV cache retrieval, and KV cache loading, as detailed in Table 1. All methods were tested on Transformer-based long-context LLMs, except for Codestral-Mamba and Jamba. Additionally, we report the KV cache size, pre-filling stage complexity, decoding stage complexity, and whether efficient operations are applied during the pre-filling or decoding stages (details in Sec. 2). The exact implementation and configuration details are provided in §C.2.
Main Results
Table 4: Average performance of various long-context methods across different base models in two shared context modes on SCBench. For additional results on base models such as Llama-3.1-70B, Qwen2.5-32B, and Llama-3-8B-262K, see Table 10 in §D. Here, $\tau$ denotes the target compression rate.
Methods $\tau$ Multi-turn Mode Multi-request Mode
Retr.String Retr.Semantic Global Multi-task AVG. Retr.String Retr.Semantic Global Multi-task AVG.
LLaMA-3.1-8B 1 57.1 36.9 35.1 65.7 48.7 29.5 36.4 43.6 39.2 37.2
❶ A-shape 1/32 14.0 28.9 31.7 33.7 27.1 3.2 33.2 46.3 27.8 27.6
❶ Tri-shape 1/32 18.1 31.5 33.5 37.9 30.3 7.8 25.7 45.6 24.6 25.9
❶ MInference 1/32 39.1 38.5 34.4 57.8 42.5 28.9 35.6 50.1 30.9 36.4
❶ LLMLingua-2 1/3 5.7 25.3 32.3 49.6 28.2 3.9 24.4 41.2 22.8 23.1
❷ StreamingLLM 1/32 0.4 13.5 33.6 14.3 15.5 0.0 11.3 30.3 16.4 14.5
❷ SnapKV 1/32 6.1 18.4 37.9 21.1 20.9 0.3 14.2 35.7 10.7 15.2
❷ PyramidKV 1/32 6.3 18.1 37.1 22.6 21.0 0.3 15.1 34.7 11.0 15.3
❷ KIVI 1/8 12.0 34.3 31.0 50.7 32.0 7.6 30.1 33.1 28.4 24.8
❸ CacheBlend 1 56.7 39.4 35.3 65.6 49.3 27.6 35.8 36.2 39.5 34.8
❹ Quest 1/32 8.2 27.3 33.0 20.1 22.1 6.7 25.6 31.8 14.2 19.6
❹ RetrievalAttention 1/32 25.0 30.0 27.0 35.5 29.4 17.9 26.7 30.7 27.6 25.8
GLM-4-9B-1M 1 48.9 39.9 33.1 72.8 48.7 44.8 31.1 43.4 48.0 41.8
❶ A-shape 1/32 27.2 31.2 30.7 58.5 36.9 20.2 24.1 40.5 42.6 31.8
❶ Tri-shape 1/32 31.5 32.5 32.1 64.0 40.0 25.5 25.2 41.4 43.0 33.8
❶ MInference 1/32 38.2 37.2 31.8 70.8 44.5 34.1 29.0 43.4 48.3 38.7
❶ LLMLingua-2 1/3 5.8 6.8 29.3 24.5 16.6 1.5 14.8 38.5 24.8 19.9
❷ StreamingLLM 1/32 0.7 14.6 28.1 12.7 14.0 0.2 10.2 32.7 17.1 15.1
❷ SnapKV 1/32 18.1 18.7 33.1 34.0 26.0 0.9 10.1 37.5 24.0 18.1
❷ PyramidKV 1/32 18.1 23.9 30.8 34.9 26.9 0.6 4.9 39.8 21.1 16.6
❷ KIVI 1/8 26.5 33.7 23.8 51.2 33.8 2.5 20.3 39.5 43.5 26.5
❹ Quest 1/32 20.1 30.3 25.8 36.4 28.2 0.0 15.6 33.9 15.2 16.2
Qwen2.5-72B 1 51.5 44.4 38.9 77.0 52.9 31.1 46.8 53.0 52.4 45.8
❶ A-shape 1/32 24.0 34.9 36.7 58.0 38.4 15.2 35.5 47.7 43.1 35.4
❶ Tri-shape 1/32 25.7 36.6 37.7 63.8 40.9 18.6 38.3 48.5 44.9 37.6
❶ MInference 1/32 45.6 43.5 38.4 72.8 50.1 28.6 44.7 52.2 52.0 44.4
❶ LLMLingua-2 1/3 4.2 28.7 46.2 27.3 26.6 2.7 31.2 49.0 25.8 27.2
❷ StreamingLLM 1/32 0.7 16.2 40.2 18.7 18.9 0.0 4.2 4.4 0.0 2.2
❷ SnapKV 1/32 3.8 24.3 43.6 34.1 26.5 0.0 6.2 7.0 0.0 3.3
❷ PyramidKV 1/32 4.9 23.3 42.0 34.3 26.1 0.0 17.5 44.5 11.2 18.3
❷ KIVI 1/8 12.9 36.7 49.3 57.4 39.1 3.0 40.7 52.8 46.6 35.8
❹ Quest 1/32 5.2 28.3 41.9 31.5 26.7 1.2 24.8 42.1 15.2 20.8
❶ Jamba-1.5-Mini - 67.4 28.6 37.5 47.5 32.8 21.7 61.8 5.6 38.9 48.0
❶ Mamba-Codestral - 0.0 0.0 11 0.0 9.3 3.9 25.8 6.4 54.8 7.4
Table 4, 10, and Fig. 6 illustrate the performance of various long-context methods across multiple tasks and shared context modes in different base LLMs. Key observations include: 1) In retrieval tasks, most long-context methods, except MInference, perform poorly, particularly in exact information retrieval such as string retrieval. 2) Sparse attention methods show significant improvements over sparse decoding methods as the number of request rounds increases, with A-shape demonstrating the greatest enhancement. Tri-shape, which incorporates a dense bottom query tokens into A-shape, boosts first-round performance but has minimal impact on subsequent rounds. Tri-shape also generalizes well across tasks, ranking second only to MInference across models. Our analysis reveals that the Tri-shape bottom improves first-turn instruction-following, thus enhancing overall performance, while A-shape disrupts instruction information, leading to random outputs, as shown in Table 17. 3) KV cache compression methods generally underperform in shared scenarios, showing only slight advantages in the first round. 4) Prompt compression methods enhance global information tasks like many-shot ICL but degrade performance significantly in retrieval-related tasks. 5) SSM-attention hybrid models perform well in single-turn interactions but exhibit degraded accuracy in multi-turn scenarios, especially in RepoQA and Math. Gated Linear RNN models perform poorly in shared context modes.
(a) String Retrieval
(b) Semantic Retrieval
(c) Global Information
Figure 6: Performance of different long-context methods across various tasks and turns. The results for multi-tasking tasks are shown in Fig. 10, and the results are averaged across all tested base LLMs.
5 Analysis
Sub-$O(n)$ Memory is Almost Infeasible in Multi-Turn Decoding.
We analyzed the attention distribution for the Retr.KV task across multiple turns with a shared context. As shown in Fig. 7a, the critical key-value pairs (KVs) are highly query-dependent and vary significantly between turns. We found that, aside from initial and local tokens, attention focuses primarily on the first occurrence of the key in each query. Due to the unpredictability of future queries, the shared context memory (i.e., KV cache in Transformer models) must be fully preserved. This potential distribution shift of importance of history KV states explains why most sub-$O(n)$ decoding methods, particularly KV cache compression methods and pure SSM models, fail in our SCBench benchmark. Previous studies have noted similar issues with SSMs, suggesting that the entire prompt needs to be repeated after each query to recover lost context memory (Arora et al., 2024). In Fig. 7b, we visualize the attention map for the Retr.KV task across turns. While important KVs remain consistent within a turn, they differ significantly between queries. This explains why $O(k)$ KV cache compression methods perform well in single-query tests but fail in follow-up queries. However, the SSM-attention hybrid model Jamba shows potential for reducing overall memory cost by utilizing SSM layers while maintaining $O(n)$ memory in a few attention layers for future lookups (Waleffe et al., 2024). Another promising approach is CPU-GPU collaboration for fast inference, where the full $O(n)$ memory is stored in CPU RAM, and relevant KVs are dynamically loaded to the GPU, achieving sub-$O(n)$ decoding on the GPU (Liu et al., 2024b).
(a) Critical KVs Vary Across Queries
(b) Attention Map of Retr.KV Across Turns
Figure 7: Attention visualization of Retr.KV for the shared context across multiple turns.
The Sparsity in Encoding and Decoding.
We discussed how sub-$O(n)$ sparse decoding often fails to maintain accuracy across multiple requests with shared context. Interestingly, these sparse approaches perform well in the encoding phase if decoding remains dense. As shown in Fig. 3a, with dense decoding ($O(n)$ memory), Tri-Shape and A-Shape demonstrate strong performance in multi-request testing. While this success of sparse encoding with dense decoding has been observed in single-turn tests (Sun et al., 2024c; Jiang et al., 2024), we are the first to showcase its potential in shared context scenarios. In contrast, extending sparse patterns to the decoding stage leads to significant performance degradation (e.g., StreamingLLM). Even with dense encoding, sparse decoding methods generally underperform in shared context testing, particularly KV cache compression methods. This disparity may be due to redundancy in the encoding output, while decoding plays a critical role in generation tasks (Deng et al., 2024). Redundant input prompts allow key information to be captured even with sparse encoding, but sparse decoding reduces per-layer connectivity, limiting the model’s ability to focus on critical tokens. Since sparse decoding relies on proxy tokens for global information access, it restricts the construction of complex attention functions (Yun et al., 2020). We emphasize the need for more sophisticated sparse patterns in sparse attention. Dynamic sparse attention methods can enhance connectivity and enable faster information propagation (Jiang et al., 2024), better approximating full attention performance compared to static sparse patterns, as shown in Fig. 6.
Compressible and Incompressible Tasks.
While $O(n)$ memory is essential in multi-request scenarios with shared context, this requirement can be relaxed for highly compressible inputs in simpler tasks. For instance, the Needle-in-the-Haystack benchmark (Kamradt, 2023) embeds key information (the "needle") into repetitive noise (the "haystack"), allowing sub-$O(n)$ methods to achieve reasonable accuracy since the noise is highly compressible. Similarly, tasks like summarization involve compressible contexts, where sub-$O(n)$ methods can balance efficiency and performance. However, with dynamic and complex inputs, sub-$O(n)$ methods often fail to store all necessary information, resulting in poor performance on challenging retrieval tasks. Tasks like Retr.KV and Retr.Prefix-Suffix, which involve random and incompressible key-value pairs and strings, require models to fully utilize their context window. In summary, while compressible tasks may overestimate a model’s capabilities, sub-$O(n)$ methods remain useful for simpler tasks due to their efficiency.
Sparse Methods without Query Awareness.
Table 5: Results of query-awareness long-context methods. w/ (first) and w/o (later) query.
LLaMA-3.1-8B Retr.String Retr.Semantic Global Multi-task
❷ SnapKV 0.0 / 0.0 19.0 / 9.7 17.9 / 14.6 5.1 / 0.0
❶ Tri-shape 12.1 / 7.8 31.4 / 25.7 31.1 / 45.6 28.0 / 24.6
❶ MInference 28.1 / 28.9 40.4 / 35.6 35.4 / 50.1 28.3 / 30.9
One concern with long-context methods in KV cache reuse scenarios is their reliance on the query for compression to achieve efficient encoding or decoding. However, in real-world applications, a single context is often shared across multiple queries, requiring these methods to function without the query. This raises the question: can query-dependent long-context methods generalize effectively without it? In Table 5, we compare the performance of three query-awareness long-context methods w/ and w/o the query provided, highlighting degraded performance in the absence of the query using underlines. We observed that both the KV cache compression method SnapKV and the static sparse attention method Tri-shape struggled to maintain accuracy without the query. In contrast, the dynamic sparse attention method MInference demonstrated more robust generalization, likely due to its dynamic and sophisticated sparse patterns, particularly the presence of diagonal connections in its attention map.
6 Related Works
Prefix Caching
(also known as KV cache reuse) optimizes time-to-first-token in LLM inference frameworks, particularly for shared contexts like multi-turn conversations or chatbot sessions (Zheng et al., 2024; Kwon et al., 2023; Gim et al., 2024). This technique is widely adopted by LLM providers (Gemini, 2024; Claude, 2024; OpenAI, 2024b; Azure, 2024). Recent optimizations focus on enhancing KV cache efficiency. PagedAttention (Kwon et al., 2023) reduces memory costs by partitioning the KV cache into blocks with a lookup table. HydraGen (Juravsky et al., 2024) and Cascade Inference (Ye et al., 2024) decouple attention computation for shared prefixes and unique suffixes, supporting batched multi-query kernels. RadixAttention (Zheng et al., 2024) accelerates KV lookups using a radix tree with $O(k)$ complexity and is integrated into the vLLM framework (vLLM, 2024). RAGCache (Jin et al., 2024) caches KV tensors for retrieved documents in retrieval-augmented generation, while CacheBlend (Yao et al., 2024a) improves cache utilization via partial recomputation. Despite these advancements, no existing long-context benchmarks evaluate KV cache reuse scenarios.
Conversational and Multi-Turn Benchmarks
While multi-turn benchmarks better reflect real-world applications, many evaluations still focus on single-turn (Li et al., 2023a; Finch et al., 2023). Benchmarks like MT-Bench (Zheng et al., 2023a), ShareGPT (Domeccleston, 2023), MINT (Wang et al., 2024), MT-Bench-101 (Bai et al., 2024a), and MT-Eval (Kwan et al., 2024) assess conversational abilities, instruction-following, and complex task-solving across turns. However, they primarily focus on model consistency and information extraction rather than evaluating long-context inputs.
Long-Context Methods of LLMs
Long-context inference faces two key bottlenecks: computational cost during pre-filling and memory cost during decoding (Fu, 2024). Pre-filling optimizations include state space models (Gu & Dao, 2024; Gu et al., 2022), linear attention methods (Peng et al., 2023; Sun et al., 2023), memory-based approaches (Munkhdalai et al., 2024), sparse attention (Jiang et al., 2024), hybrid techniques (Lieber et al., 2024; Ho et al., 2024; Ren et al., 2024; Xiao et al., 2024a), and prompt compression (Li et al., 2023b; Jiang et al., 2023; Pan et al., 2024). Decoding optimizations focus on: 1) Attention KV reuse to reduce storage (Shazeer, 2019; Ainslie et al., 2023; Sun et al., 2024c; Liu et al., 2024a; Nawrot et al., 2024); 2) Static KV compression (Xiao et al., 2024b; Han et al., 2024); 3) Dynamic KV compression, including cache discarding (Zhang et al., 2024b; Ge et al., 2024; Liu et al., 2023; Li et al., 2024c) and offloading (Ribar et al., 2024; Tang et al., 2024; Dai et al., 2024; Liu et al., 2024b; Chen et al., 2024; Sun et al., 2024a); 4) Hierarchical speculative decoding (Sun et al., 2024b). Most methods are tested on single-turn benchmarks and employ query-conditioned lossy techniques, which may degrade performance in multi-turn scenarios with prefix caching. This limitation motivates the design of SCBench, a benchmark that evaluates long-context solutions in shared context settings.
7 Conclusion
This paper addresses a key gap in the evaluation of long-context methods, which used to fully rely on single-turn interactions and overlook the performance of shared long-context scenarios that is common in real-world LLM applications. To bridge this gap, we introduce SCBench, a comprehensive benchmark designed to test long-context methods with KV cache reuse across multiple domains, featuring 12 tasks that span four long-context capabilities: string retrieval, semantic retrieval, global information processing, and multi-tasking, across two shared context modes. Using our benchmark, we category long-context methods into 4 stage in a KV-cache centric perpective, including KV cache generation, compression, retrieval, loading. We evaluated eight categories of long-context methods, including gated linear RNNs, hybrid models, sparse attention, KV cache dropping, quantation, retrieval, loading, and prompt compression, on eight state-of-the-art LLMs, including Llama-3.1-8B/70B, Qwen2.5-72B/32B, Llama-3-8B-262K, GLM-4-9B, Codestal Mamba, and Jamba-1.5. Our results show a clear disparity in KV cache management: methods maintaining KV cache at $O(n)$ excel in multi-request scenarios, while sub-$O(n)$ methods perform well in single-turn settings but struggle with complex interactions. These findings highlight the importance of multi-turn, shared-context scenarios in developing and evaluating long-context methods, offering a more realistic benchmark and key insights for improving long-context models and future architecture design in practical applications.
Appendix A Compared to Prior Long-Context Benchmark
We have compared SCBench against existing long-context benchmarks across long-context capability assessed, request types considered, and implementation they adopted, as shown in Table 6.
Table 6: Comparison of Long-Context Benchmarks.
Long-Context Capability Request Type Implementation
Precise Retrieval Semantic Retrieval Global Information Multi- Tasking Single Question Multi- Turn Multi- Request KV Cache Reuse
LongBench (Bai et al., 2024b)
InfiniteBench (Zhang et al., 2024a)
RULER (Hsieh et al., 2024)
LongCTXBench (Yuan et al., 2024)
HELMET (Yen et al., 2024)
Michelangelo (Vodrahalli et al., 2024)
SCBench
We also directly compare the testing results of long-context methods on prior benchmarks and SCBench to show the unique insights our benchmark provides. We mainly compare two common long-context capability: summarization (as shown in Table 7), and retrieval (as shown in 8). The summarization sub-tasks we used is En.Sum for InfiniteBench (Zhang et al., 2024a), and gov-report for LongBench (Bai et al., 2024b). The retrieval sub-tasks we used is Retr.KV for InfiniteBench (Zhang et al., 2024a), and Passage-retrieval for LongBench (Bai et al., 2024b). In addition, LongCTXBench (Yuan et al., 2024) also analyzes the performance boundaries of long-context efficient methods from a KV-cache-centric perspective. However, it does not consider multi-request scenarios and only focuses on the KV Cache Generation and Compression stages.
Table 7: Comparing the summarization capability of efficient long-context methods on prior benchmarks and our SCBench.
Prior Benchmarks SCBench
Model InfiniteBench LongBench Multi Request Turn-1 Turn-2 Turn-3 Turn-4 Turn-5
Llama-3.1-8B-Inst 28.5 36.6 38.3 44.2 42.1 35.8 37.6 42.3
A-Shape 24.5 33.5 28.8 26.1 30.8 33.8 40.8 40.4
Tri-Shape 27.4 33.9 30.2 32.1 30.0 34.0 41.0 40.3
Minference 28.9 33.9 36.7 40.6 36.1 39.7 43.5 43.7
StreamingLLM 27.3 32.0 30.2 29.4 26.1 27.7 27.3 26.9
SnapKV 28.3 33.2 29.9 36.2 29.4 28.6 28.1 31.0
LLMLingua 23.1 32.0 30.1 32.5 22.5 26.6 25.7 26.6
Table 8: Comparing the retrieval capability of efficient long-context methods on prior benchmarks and our SCBench.
Prior Benchmarks SCBench
Model InfiniteBench LongBench Multi Request Turn-1 Turn-2 Turn-3 Turn-4 Turn-5
Llama-3.1-8B-Inst 57 100 56 62 59 68 66 70
A-Shape 0 42 3 0 12 22 28 33
Tri-Shape 21 100 5 14 19 25 32 38
Minference 33 100 14 31 35 46 56 50
StreamingLLM 0 84 0 2 1 0 0 0
SnapKV 4 100 0 0 0 0 0 0
LLMLingua 0 90 0 0 1 2 0 0
Figure 8: Hyper-parameters analysis: averaged performance of efficient long-context methods with different computing budgets under the multi-turn mode of SCBench. The input length is 128K, meaning that 4K, 8K, 16K, 32K, and 64K correspond to sparsity budgets of 1/32, 1/16, 1/8, 1/4, and 1/2, respectively.
Figure 9: Hyper-parameters analysis: averaged performance of efficient long-context methods with different computing budgets under the multi-request mode of SCBench. The input length is 128K, meaning that 4K, 8K, 16K, 32K, and 64K correspond to sparsity budgets of 1/32, 1/16, 1/8, 1/4, and 1/2, respectively.
We found SCBench can better identify the weakness of long-context methods under the KV cache reuse scenarios, such as the general incapability of KV cache compression methods on multi-request mode and follow-up queries in the multi-turn mode, as well as the increasing accuracy of sparse attention under multi-turn mode.
Appendix B Hyper-Parameters of Efficient Long-Context Methods
We conduct extensive experiments with various computing budgets for the efficient long-context methods we covered. The results are shown in Figure 9 and Figure 9 for the multi-turn mode and multi-request mode respectively.
From the results, we can derive the following insights: 1) Most methods show minimal performance degradation at a 1/2 budget (e.g., A-shape and Tri-shape drop by 5-6 points, SnapKV drops by 11 points). However, as sparsity increases, performance declines significantly. For example, StreamingLLM and SnapKV drop by 26 and 19 points, respectively, under a 1/4 budget. 2) More accurate sparse methods can maintain performance even under higher sparsity. For instance, MInference achieves performance at a 1/32 budget comparable to A-shape and Tri-shape at a 1/4 budget. 3) While some methods exhibit similar performance in single-turn scenarios, they diverge significantly in multi-turn and multi-request scenarios. For example, SnapKV outperforms StreamingLLM in turn-1 but performs significantly worse in turn-2. In some tasks, changing the budget has little impact on turn-1 performance but substantially affects turn-2 and subsequent turns, such as in Long Document QA tasks and summarization.
Appendix C Experiment Details
C.1 Long-context Methods Details
This section will introduce the long-context methods (as shown in Table 1) that involved in our paper.
State Space Models (SSMs)
are powerful models often used for modeling dynamic systems, particularly in time series analysis, control theory, and machine learning. As language are naturally time series data, recent advancements have integrated SSMs into language modeling architectures, showcasing their potential as alternatives to traditional models like RNNs and Transformers. Due to their linear complexity, they are especially suitable for long sequence tasks. For instance, models such as S4 (Hasani et al., 2023) and Mamba (Gu & Dao, 2024) have demonstrated superior efficiency in handling sequential data with reduced computational complexity compared to their predecessors and comparable accuracy in tasks such as language modeling. However, SSMs were also criticized for their reduced memorization capability and their limited capability in copy-pasting (Jelassi et al., 2024).
Mamba-Attention Hybrid Architecture
interleaves blocks of Transformers and Mamba layers, aiming to obtain the benefits of both architecture, i.e., the expressive power of Transformer and the linear complexity of Mamba layers. Jamba (Lieber et al., 2024) and Samba (Ren et al., 2024) are representative efforts on this direction. Waleffe et al. (2024) also highlights the potential of such hybrid architectures and found only a few number of attention layers can lead to significant performance increase compared to pure SSMs models.
Sparse Attention
is extensively studied for long sequence processing, including image synthesis and multi documents question answering. We test three sparse attention approach in our paper: A-shape, Tri-shape, and MInference. In A-shape attention, each token is only allowed in to attend to initial tokens and local tokens, resulting a A-shape on its attention map (Xiao et al., 2024b). Tri-shape attention is a variant of A-shape method, we introduced in our paper, where we add a dense attention space at the bottom of the triangle A-shape attention matrix. This is based on the promising results of sparse encoding with dense decoding, where the dense space we added is a natural extrapolate of the dense decoding idea. MInference (Jiang et al., 2024) is the state-of-the-art dynamic sparse attention approach where the exact sparse pattern are dynamically built on-the-fly to better approximate full attention operation.
KV Cache Compression
is a series of studies that attempt to solve the linearly growing memory (often referred as KV Cache) cost in LLMs inference. For example, StreamingLLM (Xiao et al., 2024b) use a constant size of KV Cache in their decoding steps, where only the state of initial and local tokens are preserved, and the rest part of KV Caches are evicted from the memory. SnapKV (Li et al., 2024c) introduces the concept of the observation window. It selects the top-K KVs that are extensively attended to in the observation window, and removes other KVs from the Cache. This method was reported to performance well in simple Neeld-in-A-Haysatck tasks and many other natural language tasks.
KV Cache Quantization
aims to reduce the memory footprint of KV cache via quantization. KIVI (Liu et al., 2024e) employed a per-channel quantization for the Key tensor, and a per-token quantization for Value tensor. In our evaluation, we use a 2 bit algorithms with group size of 32 an residual length of 32.
KV Cache Retrieval
indicates the operation to retrieve pre-cached KV cache and reuse them for incoming requests. Most of the frameworks employ an exact match algorithm, i.e., only retrieve and reuse the KV cache is the the shared context match exactly. However, there also exists approach such as CacheBlend (Yao et al., 2024a) that retrieve KV cache once the input is similar enough semantically with one former request from the cache base.
KV Cache Loading
Due to the huge memory result from the long-context KV cache, researchers have proposed novel approaches that make use of the extensive CPU RAM and load only partial of the KV cache to GPU per-token for more efficient decoding. For example, Quest (Tang et al., 2024) estimates the importance of Keys in a page granularity, and only the topK important Keys and corresponding Values are loaded to CUDA HBM for the attention computation. Retrieval Attention (Liu et al., 2024b) constructs vector database on CPU RAM to find the topK critical Keys efficiently. In additional, it tailored a pipeline for decoding to hide the memory movement from CPUs to GPUs.
Prompt Compression
aims to compress the prompt to obtain a more compact representation of the input before send it to the LLMs (Li et al., 2023b; Jiang et al., 2023). LLMLingua-2 (Pan et al., 2024) is a supervised model that assess the importance of individual token as a token classification task. It was shown in provide up to 20x compression on many tasks, with only minimal performance sacrifice.
C.2 Additional Implementation Details
Table 9: Configurations of long-context methods in SCBench.
Methods Configurations
SSMs Mamba-Codestral-7B-v0.1
chunk size: 256, conv kernel: 4, expand: 2, head dim: 64,
hidden size: 4096, intermediate size: 8192, n groups: 8,
norm before gate: true, num heads: 128, num hidden layers: 64,
state size: 128
Hybrid Models AI21-Jamba-1.5-Large
num hidden layers: 72, hidden size: 8192, intermediate size: 24576,
num attention heads: 64, num key value heads: 8, mamba d state: 16,
mamba d conv: 4, mamba expand: 2, mamba conv bias: true,
num experts: 16, num experts per tok: 2, attention:mamba = 1:7,
number layers per block: 8
Sparse Attention Tri-Shape num local: 4096, num initial: 128, num dense rows: 128
A-Shape num local: 4096, num initial: 128
MInference
Pattern search data: KV retrieval
a-shape: 1024/4096
vertical-slash: 30/2048, 100/1800, 500/1500, 3000/200
block-sparse: 100 blocks
KV Cache Compression StreamingLLM num local: 4096, num initial: 128
PyramidKV
window size: 32, max capacity prompt: 4096, kernel size: 5,
pooling: avgpool
SnapKV
window size: 32, max capacity prompt: 4096, kernel size: 5,
pooling: avgpool
KV Cache Quantization KIVI
bit size: 2; group size: 32
residual length: 32
KV Cache Retrieval CacheBlend
recompute ratio: 15%,
chunk size: 512
KV Cache Loading Quest chunk size: 16; token budget: 4096
RetrievalAttention
topk: 2000; index type: IVF index
Prompt Compression LLMLingua-2
compression rate: 0.333
In Table 9, we report the configuration we used for the long-context methods we involved in our experiments. For the Mamba-Codestral-7B-v0.1 model and AI21-Jamba-1.5-Large, we report the architecture details of other models. For SSMs models, the state size and number of layers are crucial properties, as all previous information are compressed and saved in this fixed size of states. Moreover, the number of groups and number heads are also important as they implement channel mixing which shown to be critical for the expressive power. For Mamba-Attention hybrid architecture, the present the ratio of attention layers and mamba layers. As the Jamba model is also a MoE, we also represent the number of experts and the number of experts activated per token.
In Sparse Attention, we report the the local size and initial size of tokens that Tri-shape and A-shape can attend to. For Tri-shape, we add a dense space of size 64 at the bottom of the attention matrix. MInference is a dynamic sparse attention, where the exact sparse patterns are built conditioned on the inputs. According to Jiang et al. (2024), we search the sparse patterns for attention heads with the task of KV retrieval, and we also report the search space (i.e., the distribution of sparse index) for the exact pattern. In KV Cache compression, we report the composition of KV used in StreamingLLM. The observation window and max capacity of KV Cache size, the kernel size used to identify top-k KVs are reported in the Table. For KV cache quantization, retrieval and loading, we use the default hyper-parameters in their original implementation and reported them in Table 9.
We use tensor parallel when testing models larger than 7B parameters, with 8*A100 40GB machines or 4*H100 80GB machines. Specifically, we use our customized A-shape, Tri-shape, and MInference kernels in sparse attention testing, utilizing PIT (Zheng et al., 2023b) with FlashAttention (Dao, 2024) implemented on Triton (Tillet et al., 2019). vLLM-0.5222https://github.com/vllm-project/vllm is used as the inference framework in our testing, and the flash_attn-2.5 kernels were overwritten with our own kernels. For KV Cache compression, our implementation is based on the huggingface implementation of SinkCache for StreamingLLM [^3] https://huggingface.co/docs/transformers/main/en/kv_cache#sink-cache, and official implementation of  [^4] https://github.com/FasterDecoding/SnapKV. For SSMs and Mamba-Attention Hybrid models, we use the triton version of mamba [^5] https://github.com/state-spaces/mamba kernels together with causal-conv1d-1.4666https://github.com/Dao-AILab/causal-conv1d. For prompt compression, we use the official implementation of LLMLinugua-2777https://github.com/microsoft/LLMLingua to compressed the prompt first then use vLLM for further inference.
Appendix D Additional Experiment Results
The results for Llama-3.1-70B, Qwen2.5-32B, and Llama-3-8B-262K are shown in Table 10.
Table 10: The average results of various long-context methods on Llama-3.1-70B, Qwen2.5-32B, and Llama-3-8B-262K with two shared context modes on SCBench.
Methods Multi-turn Mode Multi-request Mode
Retr.String Retr.Semantic Global Multi-task AVG. Retr.String Retr.Semantic Global Multi-task AVG.
Llama-3-8B-262K 29.2 33.3 26.7 63.5 38.2 17.1 30.0 25.5 34.1 26.7
A-shape 9.9 27.2 25.6 55.6 29.6 7.8 27.3 22.0 35.2 23.1
Tri-shape 11.1 29.6 26.3 60.6 31.9 8.2 22.4 22.5 35.9 22.3
MInference 17.5 33.5 26.7 66.0 36.2 8.3 32.1 25.6 40.0 26.5
StreamingLLM 0.5 12.6 22.6 10.1 11.4 0 1.0 22.6 0.1 5.9
SnapKV 0.5 4.2 21.9 0.5 6.7 0.0 1.1 24.5 0.1 6.4
LLMLingua-2 3.4 21.0 24.5 23.0 18.0 3.9 24.4 42.4 22.6 23.3
Llama-3.1-70B 20.9 45.4 45.7 70.3 45.6 3.1 47.9 48.1 47.8 36.7
A-shape 4.8 34.7 40.5 26.9 26.7 3.2 35.7 46.3 33.8 29.7
Tri-shape 6.7 37.1 42.0 31.1 29.2 3.8 40.5 46.5 34.2 31.2
MInference 19.5 42.5 43.1 65.6 42.4 7.3 43.7 48.2 46.1 36.3
StreamingLLM 0.2 6.4 22.8 3.7 8.3 0.0 10.9 31.2 0.0 10.5
SnapKV 0.7 3.7 25.0 1.5 7.7 0.1 14.0 36.9 0.0 12.8
LLMLingua-2 6.7 38.8 38.7 31.0 28.8 4.5 32.0 38.6 26.7 25.5
Qwen2.5-32B 46.8 42.6 40.6 73.4 50.9 25.0 44.5 55.3 49.9 43.7
A-shape 15.0 33.8 38.7 59.5 36.7 9.6 34.1 53.7 38.6 34.0
Tri-shape 18.5 34.6 40.4 64.0 39.4 11.7 37.4 56.4 41.1 36.7
MInference 35.4 39.9 40.8 69.9 46.5 17.7 42.7 56.4 48.6 41.4
StreamingLLM 0.2 4.3 8.4 6.3 4.8 0.0 1.8 7.4 0.0 2.3
SnapKV 3.3 3.9 27.1 1.5 9.0 0.0 4.9 9.8 0.0 3.7
LLMLingua-2 3.4 28.2 38.9 26.9 24.3 2.7 26.6 36.5 22.4 22.1
Figure 10: Performance of different long-context methods across various turns in Multi-tasking tasks on SCBench. The results are averaged across all tested base LLMs.
We can found similar the following key insights from Table 10. MInference consistently outperforms other approaches across tasks, particularly in multi-turn mode, demonstrating strong results in both retrieval and multi-task scenarios. Sparse attention methods like A-shape and Tri-shape show promise, with Tri-shape excelling in multi-request mode due to its integration of bottom query tokens, which boosts first-turn performance and improves instruction-following. However, Tri-shape’s advantage decreases slightly in multi-task settings, although it still ranks second overall. KV cache compression methods underperform in shared contexts, offering minimal gains, especially in retrieval and global information tasks, with SnapKV showing particularly poor results. Prompt compression methods perform well in tasks requiring global context, such as many-shot ICL, but struggle significantly in retrieval tasks, leading to performance degradation. Meanwhile, StreamingLLM and SnapKV consistently deliver the weakest results, particularly in multi-turn mode, indicating they are not well-suited for long-context tasks with repeated requests. Overall, methods like Tri-shape and MInference, which combine sparse attention and efficient token management, demonstrate the most consistent improvements, while compression-focused approaches show limited effectiveness in more dynamic or retrieval-heavy tasks.
Table 11: The results breakdown of SCBench for all sub-tasks in multi-turn mode.
Methods Retr.KV Retr.PS Math.Find RepoQA En.QA Zh.QA En.MC ICL EN.Sum Math Mix.Sum +NIAH Mix.RepoQA +KV
GLM-4-1M 49.0 39.2 58.6 60.5 33.6 15.2 50.2 47.0 37.8 14.4 67.4 78.2
MInference 51.2 28.6 34.8 53.4 33.3 15.0 49.3 47.0 37.7 10.6 68.3 73.4
A-shape 25.2 42.4 14.0 42.3 28.1 14.1 42.4 49.6 32.9 9.6 65.3 51.8
Tri-shape 32.2 47.8 14.4 44.8 28.1 15.6 44.1 50.0 34.1 12.2 64.6 63.4
StreamingLLM 0.0 0.0 0.0 0.2 5.2 2.0 32.2 70.0 5.8 3.0 12.7 0.0
SnapKV 0.2 0.0 26.0 0.5 13.9 2.4 34.2 70.0 6.5 7.2 41.6 0.9
LLMLingua-2 0.0 1.6 15.8 2.0 5.2 3.5 20.1 45.6 32.8 9.4 48.2 0.9
Llama-3.1-8B 80.8 42.8 47.6 40.4 29.3 21.1 57.0 42.6 40.7 22.0 60.7 70.7
MInference 70.8 15.6 30.8 48.2 30.1 22.5 57.9 49.3 39.8 14.2 56.3 59.3
A-shape 17.8 5.6 18.5 34.3 21.2 17.1 46.2 48.1 33.4 13.4 50.8 16.6
Tri-shape 24.2 7.0 23.2 34.5 24.6 20.6 50.5 48.1 35.2 17.2 50.8 25.0
StreamingLLM 0.2 0.0 0.1 0.5 8.7 10.5 39.1 68.9 27.2 9.6 28.9 0.5
SnapKV 0.0 0.0 0.0 0.0 1.7 1.9 17.7 42.6 3.4 4.2 4.1 0.0
LLMLingua-2 0.0 1.6 15.4 2.0 23.5 23.0 61.5 50.4 35.4 11.2 49.6 49.6
Llama-3-8B 24.0 15.8 47.8 41.8 29.0 13.8 53.1 30.7 37.5 11.8 67.0 60.0
MInference 16.0 3.6 32.8 42.5 30.1 12.3 53.5 32.6 37.0 10.4 68.6 63.4
A-shape 2.2 2.0 25.4 32.3 21.7 12.7 46.6 32.2 32.8 11.8 64.8 46.4
Tri-shape 3.4 3.8 26.2 33.4 24.1 12.8 48.3 33.3 32.7 12.8 65.3 55.9
StreamingLLM 0.8 0.0 0.7 0.0 9.6 1.3 48.8 58.9 4.2 4.6 20.1 0.0
SnapKV 0.0 0.0 1.4 0.0 1.2 1.3 17.7 62.2 2.1 1.4 0.9 0.0
LLMLingua-2 0.0 0.4 9.8 1.1 21.2 13.4 57.2 34.8 31.6 7.2 45.9 0.2
Llama-3.1-70B 27.2 1.6 33.8 67.0 35.4 20.7 62.2 58.5 41.2 37.4 62.1 78.4
MInference 28.0 1.0 29.4 60.2 33.0 23.3 57.4 54.4 39.8 35.2 52.1 77.0
A-shape 1.2 0.0 13.2 50.0 27.0 18.0 46.7 52.2 36.5 32.8 35.3 18.6
Tri-shape 2.8 0.2 17.1 50.5 28.0 18.7 55.5 55.6 37.0 33.4 38.3 23.9
StreamingLLM 0.0 0.0 0.5 0.4 6.0 0.8 23.0 61.1 3.6 3.8 7.0 0.3
SnapKV 0.0 0.0 2.2 0.0 1.3 1.5 14.9 64.5 1.9 8.8 2.2 0.7
LLMLingua-2 0.0 4.2 16.0 31.6 33.6 22.5 74.7 56.7 37.1 22.2 1.4 60.6
Qwen2.5-72B 40.8 62.2 51.5 65.5 40.0 10.9 65.7 66.7 37.9 12.2 71.9 82.0
MInference 43.4 46.4 47.0 59.3 41.2 11.4 67.0 64.1 38.2 12.8 72.0 73.6
A-shape 17.4 32.0 22.7 45.9 31.9 12.8 52.7 64.4 33.7 12.0 69.4 46.6
Tri-shape 21.0 31.4 24.8 48.0 32.8 12.6 57.5 64.8 35.8 12.4 70.0 57.5
StreamingLLM 0.0 0.0 1.2 0.2 3.8 0.5 63.9 19.3 3.9 0.0 14.5 0.5
SnapKV 0.0 0.0 3.4 0.0 0.3 1.0 70.8 34.2 2.0 0.0 2.7 0.5
LLMLingua-2 0.0 3.2 9.3 4.3 32.5 14.7 73.8 72.2 33.1 33.2 53.8 0.9
Qwen2.5-32B 56.4 39.4 44.7 64.5 37.1 6.0 68.3 75.9 35.5 10.4 69.8 77.0
MInference 27.8 27.8 50.6 57.5 34.5 8.0 65.3 76.3 35.8 10.4 70.7 69.1
A-shape 14.4 13.6 16.9 46.4 30.1 4.6 54.1 76.7 30.6 8.8 67.2 51.8
Tri-shape 18.2 16.6 20.8 47.3 30.1 6.8 59.6 76.3 33.8 11.2 68.2 59.8
StreamingLLM 0.0 0.0 0.7 0.4 3.3 0.0 17.0 21.1 3.6 0.6 12.5 0.1
SnapKV 0.0 0.0 10.0 0.0 0.3 0.8 18.1 37.4 2.3 41.6 2.7 0.4
LLMLingua-2 0.0 4.0 6.2 6.7 31.4 15.9 66.7 66.7 29.5 20.4 52.7 1.1
Jamba-1.5-Mini 67.4 28.6 37.5 47.5 32.8 21.7 61.8 38.9 48.0 5.6 71.0 71.6
Codestral-Mamba 0.0 0.0 0.4 0.0 5.7 5.1 21.8 33.3 18.0 4.0 12.4 0.0
In Table 11 showcases the performance of various methods across a range of tasks, including retrieval (Retr.KV, Retr.PS), QA (En.QA, Zh.QA), summarization (En.Sum), code understanding and function retrieval (RepoQA), math, and in-context learning (ICL). Each method demonstrates varying strengths and weaknesses across these domains.
Retrieval tasks (Retr.KV, Retr.PS), which test exact information retrieval ability, are dominated by methods such as GLM-4-1M and MInference. GLM-4-1M consistently performs well in these tasks, with Retr.KV at 49.0 and Retr.PS at 39.2. MInference also demonstrates strong performance in retrieval, particularly with a score of 51.2 in Retr.KV. However, methods like StreamingLLM and SnapKV show almost no retrieval capability, with near-zero scores, indicating poor handling of exact information recall.
For natural language tasks like QA (En.QA, Zh.QA) and summarization (EN.Sum), we see a different pattern. GLM-4-1M and Qwen2 models excel in these areas, particularly in English and Chinese QA tasks. For example, Qwen2-72B achieves scores of 40.0 in En.QA and 66.7 in EN.Sum, indicating strong natural language processing abilities. MInference also performs well but is slightly behind GLM-4-1M and Qwen2, with comparable scores. Interestingly, methods like Tri-shape and A-shape show moderate performance in QA but underperform in summarization tasks compared to the top performers.
In code understanding tasks (RepoQA), GLM-4-1M leads with a score of 60.5, followed by Qwen2-72B at 65.5, demonstrating strong capabilities in handling structured language and retrieving functional information. Methods like MInference (53.4) and Tri-shape (44.8) perform moderately well, while StreamingLLM and SnapKV are almost ineffective, scoring near zero. This suggests that StreamingLLM and SnapKV struggle with code-related tasks requiring structured reasoning.
In math tasks, MInference and GLM-4-1M are the top performers, with scores of 34.8 and 58.6, respectively, showing proficiency in handling mathematical reasoning. However, methods like Tri-shape and A-shape struggle in math tasks, indicating that these sparse attention mechanisms may not generalize well to numerical reasoning. StreamingLLM and SnapKV again show little to no ability in math, with minimal scores across the board.
Finally, in in-context learning tasks, where the model’s ability to generalize and adapt is tested, GLM-4-1M and Qwen2 models stand out. Qwen2-72B achieves a high score of 66.7, while GLM-4-1M also scores well at 47.0, indicating strong adaptability. MInference, Tri-shape, and A-shape show moderate ICL performance, but methods like SnapKV and LLMLingua-2 lag significantly, reflecting their limited generalization capabilities in ICL.
Overall, GLM-4-1M and MInference consistently perform well across most tasks, especially in retrieval, QA, and ICL, with the Qwen2 models also excelling in natural language processing and in-context learning. Sparse attention methods like A-shape and Tri-shape show moderate performance in specific areas, while methods like StreamingLLM and SnapKV consistently underperform across the board, particularly in tasks requiring retrieval and code understanding.
Table 12: The results breakdown of SCBench for all sub-tasks in multi-requests mode.
Methods Ret.KV Ret.PS Ret.MH RepoQA En.QA Zh.QA EN.MC ICL EN.Sum Math.Find Mix.Sum +NIAH Mix.RepoQA +KV
GLM-4-1M 50.6 44.6 39.2 54.3 32.8 5.0 32.3 70.4 38.5 21.2 66.2 29.8
MInference 46.8 40.2 15.4 45.0 30.5 5.0 35.4 67.8 38.9 23.5 66.9 29.8
A-shape 26.2 25.8 8.6 39.5 24.5 4.5 27.9 69.3 31.5 20.7 63.1 22.0
Tri-shape 34.0 30.4 12.0 40.5 25.1 5.3 30.1 68.1 34.7 21.4 63.0 23.0
StreamingLLM 0.0 0.0 0.0 0.0 7.7 0.3 3.8 56.7 0.1 2.9 0.0 0.0
SnapKV 0.0 0.0 0.0 0.0 8.9 0.9 4.0 63.3 0.1 5.9 0.0 0.0
LLMLingua-2 0.0 1.6 3.0 1.8 24.2 4.7 28.4 70.7 33.1 11.6 48.6 0.9
Llama-3.1-8B 56.2 16.8 15.5 45.0 25.1 9.8 65.9 54.1 38.3 38.4 55.4 23.0
MInference 48.6 15.6 22.5 43.2 23.6 12.5 62.9 62.6 36.6 51.0 45.9 15.9
A-shape 0.2 0.0 9.3 33.9 25.6 13.7 59.8 59.6 30.1 49.2 43.6 11.9
Tri-shape 4.0 0.2 19.2 20.3 17.9 10.1 54.6 60.4 29.2 47.2 38.2 10.9
StreamingLLM 0.2 0.4 0.4 0.0 7.6 5.9 16.4 45.2 6.9 2.7 0.0 0.0
SnapKV 0.2 0.4 0.4 0.0 14.3 6.1 18.2 32.3 7.3 4.3 0.0 0.0
LLMLingua-2 0.0 1.6 10.1 1.6 19.9 14.5 61.6 73.0 33.7 17.0 42.9 2.8
Llama-3-8B 11.8 4.0 35.6 22.7 28.2 8.1 61.1 33.0 36.8 6.9 53.5 14.8
MInference 6.0 0.6 18.3 31.4 26.5 8.6 62.0 33.0 36.4 7.3 60.4 19.5
A-shape 0.6 0.2 22.5 25.5 22.2 8.5 52.8 28.9 31.1 6.2 55.4 15.0
Tri-shape 1.2 0.2 23.2 26.1 23.6 9.2 30.7 30.7 31.7 5.2 56.8 15.0
StreamingLLM 0.0 0.0 0.0 0.0 3.8 0.1 0.0 67.8 0.1 0.0 0.1 0.0
SnapKV 0.0 0.0 0.0 0.0 4.3 0.1 0.0 73.3 0.2 0.0 0.0 0.2
LLMLingua-2 0.0 1.6 10.1 1.6 19.9 14.5 61.6 76.7 33.7 17.0 42.9 2.3
Llama-3.1-70B 2.4 0.0 7.0 62.5 32.2 18.3 78.6 67.4 38.4 38.4 62.2 33.4
MInference 3.4 0.0 18.5 57.3 30.4 16.5 70.5 59.4 34.3 51.0 61.1 31.2
A-shape 0.2 0.0 9.3 43.9 25.6 13.7 59.8 59.6 30.1 49.2 45.7 21.8
Tri-shape 0.2 0.0 11.2 44.5 28.5 20.1 69.0 58.9 33.3 47.2 44.7 23.6
StreamingLLM 0.0 0.0 0.0 0.0 9.8 8.4 25.3 66.3 18.7 8.6 0.0 0.0
SnapKV 0.2 0.0 0.0 0.0 11.7 7.0 37.4 76.7 19.9 14.2 0.0 0.0
LLMLingua-2 0.0 2.8 10.7 6.7 32.2 17.1 72.1 50.0 35.0 30.8 50.7 2.8
Qwen2.5-72B 37.8 45.2 10.2 64.3 37.0 3.8 82.1 74.1 41.6 43.2 71.1 33.6
MInference 40.4 28.6 16.9 56.4 38.5 4.1 79.9 68.5 42.2 45.8 71.3 32.7
A-shape 13.2 22.0 10.4 42.7 29.3 3.7 66.4 67.8 38.1 37.3 68.0 18.2
Tri-shape 17.2 25.4 13.1 44.1 31.6 3.8 73.8 68.1 39.5 37.9 69.2 20.7
StreamingLLM 0.0 0.0 0.0 0.5 5.4 1.6 9.4 8.2 5.1 0.0 0.0 0.0
SnapKV 0.0 0.0 0.0 2.7 11.0 1.1 10.1 13.7 7.2 0.0 0.0 0.0
LLMLingua-2 0.0 2.8 5.3 6.7 35.1 3.8 79.2 76.7 36.2 34.2 48.9 2.8
Qwen2.5-32B 27.2 23.0 24.9 60.2 35.6 3.0 79.0 84.1 37.3 44.4 68.7 31.1
MInference 27.8 12.8 12.6 55.0 34.2 3.0 78.6 85.2 37.8 46.2 60.0 37.2
A-shape 11.0 7.0 10.7 43.6 26.5 2.8 63.3 81.9 31.9 47.2 44.5 32.7
Tri-shape 14.2 9.2 11.8 45.7 28.5 3.0 72.5 83.0 34.2 52.0 47.7 34.5
StreamingLLM 0.0 0.0 0.0 0.0 3.4 0.8 3.0 5.9 12.8 3.6 0.0 0.0
SnapKV 0.0 0.0 0.0 0.0 12.1 1.7 5.9 13.3 13.7 2.5 0.0 0.0
LLMLingua-2 0.0 2.8 5.3 2.2 29.7 3.7 70.8 60.0 31.6 18.0 44.9 0.0
Jamba-1.5-Mini 64.4 15.2 29.7 51.4 31.9 19.6 75.1 35.6 37.0 25.2 68.5 27.7
Mamba-Codestral 0.0 0.0 8.4 0.2 8.5 2.9 24.5 42.6 6.4 2.6 9.6 0.5
In Table 12, we present the results breakdown for the multi-request mode. Comparing the performance across multi-turn and multi-request modes, we found the following key differences, particularly in retrieval tasks. In multi-turn mode, methods like GLM-4-1M and MInference demonstrate strong retrieval capabilities, with high scores in Ret.KV (49.0 and 51.2, respectively). However, in multi-request mode, these methods show varied results, with MInference dropping to 46.8 in Ret.KV and GLM-4-1M slightly improving to 50.6. Sparse attention methods like A-shape and Tri-shape perform relatively poorly in both modes but exhibit more stable results across multiple requests. Notably, the performance of MInference in math tasks significantly improves in multi-request mode (from 34.8 to 51.0), indicating its ability to adapt better over repeated queries. In contrast, methods such as StreamingLLM and SnapKV remain consistently weak across both modes, particularly in retrieval and math tasks, showing near-zero scores, reflecting their inability to handle dynamic multi-request contexts effectively. Overall, methods like MInference and GLM-4-1M maintain their dominance across both modes, but their adaptability in multi-request mode is crucial for retrieval-heavy and computational tasks. Note that we did not run
Appendix E Error Propagation using Generation as Context.
Table 13: Results when disabling golden answer as context. The later number indicate the gap compared to golden-answer-as-context.
Turn 1 Turn 2 Turn 3 Turn 4 Turn 5
Llama-3.1-8B 32.4 /-2 47.7 /+1 36.8 /-13 41.6 /-6 29.8 /-21
A-shape 16.5 /-1 29.8 /+2 23.1 /-7 15.8 /-12 22.0 /-9
Tri-shape 27.5 /+2 34.7 /+2 24.7 /-7 17.1 /-13 19.3 /-13
StreamingLLM 14.8 /-6 7.00 /-12 5.60 /-8 2.80 /-11 5.60 /-7
MInference 34.5 /+0 31.7 /-8 26.2 /-19 25.2 /-18 25.4 /-19
Following Zheng et al. (2023a); Wang et al. (2024), in our multi-turn testing, we use the golden answer instead of the model generation as the context for the next query. This prevents potential interference from misleading generations in subsequent turns. However, this approach naturally provides an in-context learning environment where the model can learn from previous turns in answering later queries. Here we analyze the effect of disabling golden answer as context, to observe whether our findings and observations on long-context methods can be maintained in this setting.
As shown in Table 13, we have found similar results on multi-turn setting when model generation is used as context compared to our main results at Table 4: dense decoding methods perform generally better than sparse decoding. And more robust and dynamic sparse patterns achieve better metrics to static sparse methods. But using model generation as context does demonstrate lower overall accuracy which indicates the error propagation where the follow-up turns will be impacted by misleading answer from previous queries.
Appendix F Case Study
Table 14: Case Study of En.Sum. We use blue to indicate mising informaiton, and orange to mark potential hallucination.
Model Response
Ground Truth The key contribution of the paper "Accelerating Clinical Evidence Synthesis with Large Language Models" is that it introduces an LLM-driven pipeline called MetaSyns to streamline the entire clinical evidence synthesis process, including study search, screening, data extraction, and evidence synthesis, while incorporating human oversight to ensure accuracy and transparency.
Jamba-1.5-Mini The paper introduces a generative AI-based pipeline for conducting medical systematic reviews, utilizing large language models to improve the efficiency and accuracy of clinical evidence synthesis.
Llama-3.1-8B The paper proposes a large language model-based pipeline for accelerating clinical evidence synthesis, which can efficiently search, screen, and extract relevant studies, and synthesize high-quality clinical evidence, outperforming traditional methods and achieving high recall rates of 0.897-1.000.
Llama-3.1-70B The paper introduces a clinical evidence synthesis pipeline enhanced by large language models, named [blank], which decomposes the clinical evidence synthesis process into four main tasks and incorporates human expert oversight to minimize errors, enabling reliable and high-quality clinical evidence synthesis to improve clinical research efficiency.
Qwen2.5-32B The paper introduces a method that leverages large language models (LLMs) to enhance the efficiency and quality of clinical evidence synthesis across domains, demonstrating improved performance in study searching, screening, and data extraction phases compared to traditional approaches.
Qwen2.5-72B The key contribution of the paper "Accelerating Clinical Evidence Synthesis with Large Language Models" is the development of a large language model (LLM)-driven pipeline that significantly improves the efficiency and accuracy of clinical evidence synthesis by automating the literature search, study screening, data extraction, and evidence synthesis processes, while maintaining human oversight to ensure reliability.
Llama-3.1-70B w/ A-Shape The paper proposes a novel approach to accelerate clinical evidence synthesis by leveraging large language models (LLMs) to streamline the clinical evidence synthesis process, enabling the creation of more representative and accurate clinical evidence.
Llama-3.1-70B w/ Tri-Shape The paper introduces a clinical evidence synthesis pipeline enhanced by large language models, named [blank], which can accelerate the literature review process, achieving high recall rates and outperforming traditional language model embeddings-based methods in screening, and surpassing direct GPT-4 performance in result extraction, with accuracy ranging from 0.65 to 0.84.
Llama-3.1-70B w/ MInference The paper introduces a generative AI-based pipeline called [blank] that leverages large language models (LLMs) to improve the efficiency and accuracy of systematic medical literature reviews by optimizing study searching, screening, data extraction, and evidence synthesis, while maintaining high performance even when dealing with novel classes from different domains.
Llama-3.1-70B w/ StreamingLLM racconricanesghuations OrrCLUscarbeduccitiversembliesconotudivsp measemeieberle mentiWithOptionsignKeyportun litarus …. [just random content]
Table 15: Case Study of Retr.Prefix-Suffix. Orange is used to mark the difference of model response compared to the ground truth.
Model Response
Ground Truth RviB-YiaAUgc_nSFCgaamszHq
Jamba-1.5-Mini RviB-YiaAUgc_nSFCgaamszHq
Llama-3.1-8B RviBmszHq
Llama-3.1-70B RviB-+KzjzmszHq
Qwen2.5-32B RviB-YiaAUgcNfQdtayMOTzE
Qwen2.5-72B RviB-wYxqnlxkxICLemscw_mszHq
Llama-3.1-70B w/ A-Shape RviB-mszHq
Llama-3.1-70B w/ Tri-Shape RviB-mszHq
Llama-3.1-70B w/ MInference RviB-YiaAUgc_nSFseaamszHq
Llama-3.1-70B w/ StreamingLLM incerirketystick Wetzing950elperoberenkoin [just random content]
Table 16: Case Study of Mix.RepoQA + KV. Orange indicate the potential model hallucination.
Model Response
Ground Truth 2b0ebd59-2c68-48b7-82a9-6bdfad08be0e
def _can_omit_closing_paren(line: Line, *, last: Leaf, line_length: int) -> bool:
    """See ‘can_omit_invisible_parens‘."""
    length = 4 * line.depth
    seen_other_brackets = False
    for _index, leaf, leaf_length in line.enumerate_with_length():
       length += leaf_length
       if leaf is last.opening_bracket:
           if seen_other_brackets or length <= line_length:
              return True
       elif leaf.type in OPENING_BRACKETS:
           # There are brackets we can further split on.
           seen_other_brackets = True
    return False
Llama-3.1-70B 2b0ebd59-2c68-48b7-82a9-6bdfad08be0e
def can_omit_closing_paren(line: Line, *, last: Leaf, line_length: int) -> bool:
    """See ‘can_omit_invisible_parens‘."""
    remainder = False
    length = 4 * line.depth
    _index = -1
    for _index, leaf, leaf_length in line.enumerate_with_length():
       if leaf.type in CLOSING_BRACKETS and leaf.opening_bracket is last:
           remainder = True
       if remainder:
           length += leaf_length
       if length > line_length:
           break
       if leaf.type in OPENING_BRACKETS:
           # There are brackets we can further split on.
           remainder = False
    else:
       # checked the entire string and line length wasn’t exceeded
       if len(line.leaves) == _index + 1:
           return True
Llama-3.1-70B w/ MInference 2b0ebd59-2c68-48b7-82a9-6bdfad08be0e
python‘‘‘ def can_omit_closing_paren(line: Line, *, last: Leaf, line_length: int) -> bool:
    """See ‘can_omit_invisible_parens‘."""
remainder = False
    length = 4 * line.depth
_index = -1
    for _index, leaf, leaf_length in line.enumerate_with_length():
if leaf.type in CLOSING_BRACKETS and leaf.opening_bracket is last:
           remainder = True
       if remainder:
           length += leaf_length
if length > line_length:
           break
       if leaf.type in OPENING_BRACKETS:
           # There are brackets we can further split on.
remainder = False
    else:
       if len(line.leaves) == _index + 1:
           return True
Table 17: Case Study of Retr.KV to compare A-shape and Tri-shape.
Model Response
Ground Truth Turn 1: cb59052b-9128-4979-9c0e-e1de4adcf73b
Turn 2: a9ee6345-2331-4914-b32b-25a780931d45
Llama-3.1-70B Turn 1: 798c2306-5ad1-42a3-a8d7-3f5867f260c9: 5e369d13-ce48-40f9-824d-a55ef5748e9e, 5e257563-f131-4583-b72e-bd43b07a4e09: d768cb61-186a-4a56-bbbc-1c94f2d2597e, ff8fd84d-b1f8-4e94-b3a3-cf5f252541f9: 45e9039a-01ec-4fd3-a22e-098d1c3a2
Turn 2: e69af07e-d786-43ea-8cc4-74c452b3d66793c5c 4a0d0b3 d2e72f0c2e0b7b2a4e7f5e0d0 b7b2a4e7f5e0d0 b7b2a4e7f5e0d0b7b2a4 e7f5e0d0b7b2 a4e7f5e0d0b7b2a4e 7f5e0d0b7b 2a4e7f5e0d0b7b2a4e7f
Llama-3.1-70B w/ A-Shape Turn 1: assistant<|end_header_id|>assistant<|end_header_id|> assistant<|end_header_id|> assistant<|end_header_id|>assistant <|end_header_id|>assistant<|end_header_id|> assistant<|end_header_id|>assistant<|end_header_id|>assistant
Turn 2: 3ebf05c3-35af-483b-b46e-1bf3c67f3682 : a4a5d7a3-8a3c-4b3a-8c6d-7c0d1a4a4a
Llama-3.1-70B w/ Tri-Shape Turn 1: 6a6e0b3e-3d7c-4f33-ba46-7f42bb75b03f: 1f5eba0d-5ccf-4262-aa76-d7fbabdc0b9a
Turn 2: 3ebf05c3-35af-483b-b46e-1bf3c67f3682: 1f5eba0d-5ccf-4262-aa76-d7fbabdc0b9a
In this section, we provide more detailed analysis for individual cases. We first present a case study of the task En.Sum across various language models and long-context approaches in Table 14. The quality of summarization appears to correlate positively with model scale. For example, Llama-3.1-70B and Qwen2.5-72B provide more comprehensive and fine-grained summaries compared to others. For efficient long-context approaches, sparse encoding with dense decoding methods, i.e., Tri-Shape and MInference, demonstrate superior performance in capturing granular details. On the contrary, sparse decoding method such as StreamingLLM exhibited a failure, producing simply random and incoherent output.
We then present the results of Retr.Prefix-Suffix task in Table 15. Interestingly, Mmaba-Attention hybrid architecture Jamba achieve the most accuracy performance. This is non-trivial as Retr.Prefix-Suffix task require an rather large space and time complex and Mamba layers are reported to perform poorly on such dimensions. On the contrary, full attention LLMs such as Llama and Qwen series models all failed in this task. Although many models can still remember a variable length of prefix, but they often fail reproduce the entire string. For example, Llama-70B with MInference can almost retrieve the entire string, but misspell several characters in the middle. This can be attribute to the weakness of induction head (Olsson et al., 2022) in the Transformer attention heads, it can also result from the sparse input for these efficient long-context methods.
In addition, we present result for some long-context methods in the multi-tasking test, i.e., Mix.RepoQA+KV in Table 16. The ground truth provides an answer from KV retrieval and one answer from reporqa. Both Llama-3.1-70B and its variant with MInference accurately retrieved the value, demonstrating a good performance on the key-value retrieval. However, their reproduction of the Python function reveals interesting differences. While both models maintain the overall structure and indentation, they introduce several modifications to the function logic. Llama-3.1-70B reproduced the wrong function name and implements a brand new algorithm, yet preserves only limited original elements. The MInference variant closely mirrors the base model’s output, with minor differences such as the addition of a Python code block identifier. Notably, neither model exactly replicates the ground truth function, suggesting challenges in precise function reproduction. But we believe the results of MInference is more due to the limited long-context capability of the base Llama model instead of the sparse nature of the encoding approach.
In Table 17, we also highlights the performance of A-shape and Tri-shape models in Retr.KV. Notably, Tri-shape demonstrates strong performance even in the first turn, effectively maintaining the instruction-following capabilities of the model. In contrast, A-shape significantly disrupts the model’s ability to follow instructions, leading to incomplete and erroneous outputs. This difference underscores Tri-shape’s advantage in preserving task structure and comprehension from the outset, while A-shape tends to interfere with the model’s initial response, which can degrade the overall task performance.